OpenStudy (anonymous):

You started on a bike trip, riding at a rate of 9 miles per hour. Your friend left one hour later and rode at the rate of 12 miles per hour. How long did it take your friend to catch up with you?

OpenStudy (anonymous):

You're 9 miles ahead when the 2nd bike starts. His speed wrt you is 3 mph (12-9). 9/3 = 3 hours

OpenStudy (anonymous):

So it takes 3hours

OpenStudy (anonymous):

Medal please.

OpenStudy (anonymous):

I don't understand this.. I saw that same explanation on algebra.com and it doesn't make sense to me.

OpenStudy (anonymous):

okay so your friend named oscar the grouch starts of an hour later, right?

OpenStudy (anonymous):

by that time you sarah has covered 9 miles

OpenStudy (anonymous):

but oscar is 3 miles faster ...

OpenStudy (anonymous):

I also have to find the error in someone else's work for this problem. Her work is: distance=rate x time. your distance= your friends distance. let t= your time(in hours) let t+1=your friends time (in hours). 9t=12(t+1) 9t=12t+12 -3t=12 t=-4 A negative time does not make sense in this context, so the problem has no solution.

OpenStudy (anonymous):

So to take the lead he needs to cover 9 miles

OpenStudy (anonymous):

so 9/3 =3 voila

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

yes?no?maybe?i dont know?can you repeat?

OpenStudy (anonymous):

I get it but I don't understand how to do the work.

OpenStudy (anonymous):

okay, so use the equation 9+9x=12x

OpenStudy (anonymous):

@sarahruth ?

OpenStudy (anonymous):

so would it be 1.5?

OpenStudy (anonymous):

how did you get 1.5?

OpenStudy (anonymous):

9+9x=12x 9=3x 3=x